OpenStudy (anonymous):

the plane through (-1 , 2 , -5) that is perpendicular to the planes 2x-y+z=1 and x+y-2z=3.

OpenStudy (anonymous):

@ganeshie8 please solve this too .

ganeshie8 (ganeshie8):

cross the normals of given planes, you will get the normal of required plane

ganeshie8 (ganeshie8):

normal of required plane = \(\left<2, -1, 1\right> \times \left<1, 1, -2\right>\)

ganeshie8 (ganeshie8):

since u knw a point on plance already, you're done.

OpenStudy (anonymous):

okay thanks :)

ganeshie8 (ganeshie8):

u wlc !